### v2

# augmented reality AP average precision

AR
augmented reality
AP
average precision
API
application programming interface
BEV
bird eye view
BRB
Bottleneck residual block
mBIoU
mean Boundary Intersection over Union
POI
Point of Interest
CAI
computer-assisted intervention
CE
cross entropy
CAD
computer-aided design
CNN
convolutional neural network
CRF
conditional random fields
DPC
dense prediction cells
DLA
deep layer aggregation
DNN
deep neural network
DA
domain adaption
DR
domain randomization
FAT
falling things
FCN
fully convolutional network
FoV
field of view
FV
front view
FP
False Positive
FPN
feature Pyramid network
FN
False Negative
FMSS
fast motion sickness scale
GAN
generative adversarial network
GCN
graph convolutional network
GS
Gaussian Splatting
$G$
Gaussian Grouping
HMI
Human-Machine-Interaction
HMD
Head Mounted Display
IoU
intersection over union
IRB
inverted residual bock
IPQ
igroup presence questionnaire
KNN
k-nearest-neighbor
LiDAR
light detection and ranging
LSFE
large scale feature extractor
LLM
large language model
mAP
mean average precision
MC
mismatch correction module
mIoU
mean intersection over union
MIS
Minimally Invasive Surgery
MSDL
Multi-Scale Dice Loss
ML
Machine Learning
MLP
multilayer perception
mIoU
mean Intersection over Union
NN
neural network
NDDS
NVIDIA Deep Learning Data Synthesizer
NOCS
Normalized Object Coordiante Space
NeRF
Neural Radiance Fields
NVISII
NVIDIA Scene Imaging Interface
NGP
Neural Graphics Primitives
OR
Operating Room
PBR
physically based rendering
PSNR
peak signal-to-noise ratio
PnP
Perspective-n-Point
RV
range view
RoI
region of interest
$B$
residual basic block
RAS
robot-assisted surgery
SPP
spatial pyramid pooling
SH
spherical harmonics
SGD
stochastic gradient descent
SDF
signed distance field
SfM
structure-from-motion
SAM
Segment-Anything
SUS
system usability scale
SSIM
structural similarity index measure
SfM
structure from motion
SLAM
simultaneous localization and mapping
TP
True Positive
TN
True Negative
thor
The House Of inteRactions
TSDF
truncated signed distance function
VR
Virtual Reality
YCB
Yale-CMU-Berkeley
AR
augmented reality
ATE
absolute trajectory error
BVIP
blind or visually impaired people
CNN
convolutional neural network
c2f
coarse-to-fine
FoV
field of view
GAN
generative adversarial network
GCN
graph convolutional Network
GNN
Graph Neural Network
HMI
Human-Machine-Interaction
HMD
head-mounted display
MR
mixed reality
IoT
internet of things
LLFF
Local Light Field Fusion
BLEFF
Blender Forward Facing
LPIPS
learned perceptual image patch similarity
NeRF
neural radiance fields
NVS
novel view synthesis
MLP
multilayer perceptron
MRS
Mixed Region Sampling
OR
Operating Room
PBR
physically based rendering
PSNR
peak signal-to-noise ratio
PnP
Perspective-n-Point
SUS
system usability scale
SSIM
similarity index measure
SfM
structure from motion
SLAM
simultaneous localization and mapping
TP
True Positive
TN
True Negative
thor
The House Of inteRactions
UEQ
User Experience Questionnaire
VR
virtual reality
WHO
World Health Organization
XR
extended reality
YCB
Yale-CMU-Berkeley
YOLO
you only look once
CoRe-GS: Coarse-to-Refined Gaussian Splatting
with Semantic Object Focus
Hannah Schieber1, Dominik Frischmann1, Victor Schaack1, Simon Boche3, Angela Schoellig3,
Stefan Leutenegger3 and Daniel Roth1
1Hannah Schieber, Dominik Frischmann, Victor Schaack, and Daniel Roth are with the Technical University of Munich, Human-Centered Computing and Extended Reality Lab, TUM University Hospital, Clinic for Orthopedics and Sports Orthopedics, Munich Institute of Robotics and Machine Intelligence (MIRMI) hannah.schieber@tum.de, daniel.roth@tum.de2Simon Boche is with Technical University of Munich simon.boche@tum.de3Angela Schoellig is with Technical University of Munich angela.schoellig@tum.de2Stefan Leutenegger is with the Mobile Robotics Lab, Department of Mechanical and Process Engineering, ETH Zurich lestefan@ethz.ch
Abstract
Mobile reconstruction has the potential to support time-critical tasks such as tele-guidance and disaster response, where operators must quickly gain an accurate understanding of the environment. Full high-fidelity scene reconstruction is computationally expensive and often unnecessary when only specific points of interest (POIs) matter for timely decision making. We address this challenge with CoRe-GS, a semantic POI-focused extension of Gaussian Splatting (GS). Instead of optimizing every scene element uniformly, CoRe-GS first produces a fast segmentation-ready GS representation and then selectively refines splats belonging to semantically relevant POIs detected during data acquisition. This targeted refinement reduces training time to 25% compared to full semantic GS while improving novel view synthesis quality in the areas that matter most. We validate CoRe-GS on both real-world (SCRREAM) and synthetic (NeRDS 360) datasets, demonstrating that prioritizing POIs enables faster and higher-quality mobile reconstruction tailored to operational needs.
I INTRODUCTION
Capturing and reconstructing real-world environments in 3D is increasingly essential for domains ranging from aerial imaging and disaster response to medicine, cultural heritage, and robotic navigation [1, 2, 3, 4, 5, 6]. 3D representations, like Gaussian Splatting (GS) [7] provide detailed, explorable models that support simulation, remote mission planning and operator decision making, especially in drone and robot captures. In disaster response, rapid situational awareness is critical because teams often need actionable 3D maps within minutes rather than hours. Full-scene reconstruction can be computationally costly, time-consuming, and unnecessary when only specific POI are operationally relevant. Focusing reconstruction on mission-critical regions rather than entire environments cuts processing time and memory requirements and enables faster deployment and decision-making in time-sensitive operations.
To further optimize such pipelines, especially in drone-based visual captures, Point of Interest (POI) selection can be a valuable strategy to focus on relevant areas. During capture, a POI typically guides the camera towards a specific object or region. When constructing 3D representations from these images, POI metadata can continue to play a role, informing downstream semantic segmentation tasks, guiding training prioritization, or supporting scene understanding. While learning-based 3D representations of the entire scene often require significant training overhead, integrating POI awareness can streamline this process. Currently, semantic approaches allowing the extraction of POI objects often suffer from outliers (floaters), see Fig. 1.
Figure 1: Comparison of traditional semantic GS following Ye et al. [8] compared to our CoRe-GS leveraging directly the knowledge of specific POIs during training time.
For scene refinement, active view selection [9, 10] is the standard method. However, these approaches typically optimize the entire scene. Optimizing only specific parts is usually not done, and post-processing approaches suffer from outliers (floaters) [4, 11, 8]. Semantic segmentation can be used to optimize particular parts, but incorporating these often introduces computational overhead. Although semantic GS offers sufficient fidelity to support selective refinement, it is generally applied to the entire scene [4, 8, 12]. This extends the overall runtime for initial reconstruction.
To address this challenge, we propose CoRe-GS, a refinement of POIs in GS. CoRe-GS reduces training time for high-quality scene reconstruction by focusing on a single POI. We extend classical and semantic GS by integrating POI extraction, accelerating the overall training process. Once a segmentation-ready scene is created, further processing refines only the extracted object, the POI, in roughly a quarter of the time required by standard semantic GS [8]. By effectively mitigating color space discrepancies during POI refinement, our approach enhances the visual quality of the resulting scenes. We demonstrate our method in both indoor and outdoor settings using scenes from the SCRREAM [13] and NeRDS 360 [14] datasets.
To summarize, we contribute:
  • A novel POI refinement for semantic object focus and efficient editable GS.
  • A novel color-based effective filtering method, for GS POI refinement, addressing outliers.
  • A comprehensive evaluation on diverse datasets, testing our approach on both indoor and outdoor scenes with different POI sizes to assess adaptability.
II RELATED WORK
II-A 3D Scene Refinement
Various strategies exist for scene refinement. A naive but straightforward approach is to capture more images, while more advanced methods target selective views [10], path trajectory planning [15, 16], or a combination of both [9].
While enhancing the visual quality of the scene representation is one approach, Peralta et al. [15] focus on adapting the capturing path. Similarly, Chen et al. [16] introduce GenNBV, which optimizes drone camera paths using reinforcement learning. While these approaches apply to a broader scope, others directly target GS optimization [10]. FisherRF [10] applies uncertainty quantification to select ideal views within GS. Li et al. [9] leverage efficient viewpoint sampling and path planning combined with a dense scene representation via GS.
Refinement of the scene based on semantic POIs improves efficiency in another direction than optimizing the initial capturing path. With POIs-based refinement, only the relevant parts are updated leading to a runtime improvement.
II-B Editable Radiance Fields
Editable radiance fields often employ semantics or language guidance to adapt the final scene representation [12, 8, 4, 17, 18, 19]. Different applications can be considered, ranging from object removal /extraction to object search or inpainting.
II-B1 Object Removal
Although GS [7] now dominates radiance fields, its predecessor neural radiance fields (NeRF[20, 21, 22] already offered editable approaches [21, 22]. Inpainting can remove objects from the learned radiance field using an inpainting training step [23, 24, 25]. Yin et al. [23] leveraged SAM [26] and LaMa [27] to inpaint selected objectsfor NeRF. In a similar approach, Weder et al. [24] proposed an inpainting-based NeRF solution to remove unwanted objects. Similarly, object removal tasks have been addressed in GS using segmentation masks, for example, Huang et al. [25], address this with SAM and depth-guided inpainting. In addition, semantic GS approaches provide another way to handle object removal [12, 8] and, when combined with an inpainting model such as LaMa [27], enable inpainting of Gaussians.
II-B2 Semantic Search and Editing
Although NeRF already showed great potential, GS [7] offers better rendering capability and adaptability due to its explicit scene representation. Often, multiple models are combined. Yu et al. [18] leverage CLIP, Droid-SLAM [28], and GS for robotic navigation. In addition to language guidance, HAMMER [29] produces semantic GS maps from multiple input devices, such as robots and smart glasses, allowing for direct editing of the resulting scene representations.
Similarly, Qu et al. [30] proposed a point-based editing approach that enables the manipulation of Gaussians directly within the scene. Furthermore, segmentation-based GS allows scene editing per class, often proposed as a downstream task [8, 12]. Ye et al. [8] proposed segmentation-based grouping, with the so-called Gaussian Grouping (GG), extending the Gaussian representation with a feature field for classifier-based segmentation. Moreover, they demonstrated scene editing, such as inpainting and object removal, as a downstream task. Within GAGA, Lyu et al. [12] proposed segmentation-based GS using a memory bank. Although segmentation quality can be improved with GAGA, training time also increases, as the approach uses pure GS [7] as input, followed by several semantic training and label preparation steps. For large objects and editing large scenes, Schieber et al. [4] show that a supervised model can remove objects more effectively compared to foundation models. Although this is promising for large objects, foundation models offer broader scalability for unknown classes.
II-C Applied Editable Radiance Fields
Adding semantic segmentation to the radiance field representation provides the opportunity to later edit the explicit GS scene representation. Semantic GS approaches [12, 8, 4] demonstrate this ability in a post-processing step. Others directly integrate it into robotic applications [31, 32]. Li et al. [32] added semantics for robotic manipulation for a limited number of views. They address language-guided grasping. Another grasping solution is proposed by Ji et al. [31]. They leverage GS, MobileSAM, and CLIP to propose potential object grasp positions.
II-D Research Gap
Scene editing [8, 4, 19], path planning [15, 16], and camera selection [10] have already been extensively researched, but there is still room for improvement in optimizing the semantic representations used in these contexts. Complete semantic training is costly, and subsequent POI extraction often leads to inconsistencies (e.g., floaters). We introduce CoRe-GS, a POI refinement approach GS that uses a segmentation-ready GS for initial understanding and selectively refines POI, achieving high-quality results with significantly reduced training time.
Figure 2: Our approach first trains GG [8] (initial segmentation-ready GS) for 3k initial iterations. Followed by that, the POI is extracted based on the given segmentation class (POI selection). In a last step, the POI is refined using our periodic scene filtering with the previously extracted furthest color (mask and color extraction and POI refinement).
III METHOD
III-A Gaussian Segmentation and Point-of-Interest Extraction
We apply the training process of semantic GS [8] for 3k iterations and obtain a coarse, segmentation-ready scene representation, see Figure 2. The pre-segmentation step enables each part of the scene to be represented by a semantic class or instance class, depending on the given database. During pre-extraction, POIs are selected by class ID (e.g., “car”, see Fig. 2). Images containing the chosen class are retained, and only Gaussians predominantly associated with the selected POI are preserved.
For each retained image, binary masks ($M$) are generated to isolate the target object: for each pixel ($u,v$) corresponding to the segmentation class, a value of $1$ is assigned, while all other pixels are set to 0. These binary masks ($M$) are used to effectively remove additional information. This ensures, in the later refinement process, that only the relevant parts of the scene within the POI are processed in the refinement step. During training, GS often produces floaters, unwanted Gaussians, in the background. Relying solely on a masked loss during refinement leads to soft or fuzzy contours, as areas outside the POI are effectively ignored. Alternatively, replacing the background with a generic color like black or white can result in floaters adopting those common colors, which are also likely to appear in the foreground object. This makes it difficult to distinguish and filter them later. To address both issues, we introduce a color filtering step that assigns a background color maximally distinct from the scene content, enabling effective removal of floaters based on color distance.
To employ it, we analyze the color distribution of the image. For each image, we determine the most distant color from the colors present in the image. We speed up this process by downscaling the images by a factor of $0.5$. Unique furthest colors are then extracted from the downscaled images, and a KD-tree is created based on these furthest colors. The set of existing image colors is given by:
$C=\{c_{i}\}_{i=1}^{N},\quad\text{where}\quad c_{i}\in[0,1]^{3}\in\rm I\!R$ (1)
A predefined reduced RGB color space ($P$):
$P=\{p_{j}\}_{j=1}^{J},\quad\text{where}\quad p_{j}\in[0,1]^{3}\in\rm I\!R$ (2)
is sampled systematically across the RGB spectrum ($r,g,b$). $P$ is queried against the constructed KD-tree. The color with the greatest minimum Euclidean distance ($\|p_{j}-c_{i}\|_{2}$) from all existing image colors is identified as the “furthest” color $p^{*}$, and later used for rendering, to color floaters uniquely:
$p^{*}=\arg\max_{p_{j}\in P}\bigg{(}\min_{c_{i}\in C}\|p_{j}-c_{i}\|_{2}\bigg{)}$ (3)
where the Euclidean distance in RGB space is given by:
$\|p_{j}-c_{i}\|_{2}=\sqrt{(r_{p_{j}}-r_{c_{i}})^{2}+(g_{p_{j}}-g_{c_{i}})^{2}+(b_{p_{j}}-b_{c_{i}})^{2}}$ (4)
where $r,g$, and $b$ represent the RGB values normalized to the interval $[0,1]$. This information is utilized during the refinement process to render the background color and to identify unwanted background floaters.
Figure 3: Exemplary occlusion removal comparison on the scene SF_6thAndMission_medium6 (bottom) and SF_6thAndMission_medium10 (top) after 30,000 iterations. Post-training isolation using $G$ partially recovers occlusions, with “convex hull” usage influencing the result. Our approach produces a clean background.
III-B Point of Interest Refinement
The “furthest” color is used as a universal background for the subsequent rendering. To focus refinement on the selected POI, the training images and loss calculations are masked with binary masks $M(x,y)$. We initialize our POI refinement in CoRe-GS using the filtered points of the POI and start the training process for the remaining 27,000 iterations. In addition, our GS is regularly filtered using the “furthest” color to eliminate potential floaters during refinement.
III-B1 Periodic Scene Filtering
Although segmentation classes from a 3D semantic GS can initially isolate the POI, they are not used during refinement because the semantic rasterizer and classifier significantly increase training time. Instead, we employ our color-based filtering strategy for additional pruning and adjustment of the POI GS. In the pre-extraction step, we received the most distinct (“furthest”) color ($p^{*}$) relative to the colors already present, measured by Euclidean distance in RGB color space ($d_{avg}$).
Using $p^{*}$, we effectively filter out unwanted Gaussians. Given the substantial color difference between $p^{*}$ and meaningful Gaussians, unwanted artifacts (floaters) can be identified and removed through color-based filtering. Specifically, for each view, the color of each Gaussian is computed by evaluating its spherical harmonics (SH) coefficients relative to the camera’s viewing direction. The resulting RGB color is then compared to the previously identified ($p^{*}$). Then the Euclidean distance between the Gaussians computed RGB color and $p^{*}$ is calculated. If this distance falls below a threshold, the Gaussian is considered an artifact and marked for removal. All Gaussians identified as artifacts based on this criterion are pruned from the scene representation, thereby significantly reducing unwanted visual artifacts that typically manifest around object boundaries. Given the average distance ($d_{avg}$) computed earlier from selecting $P$, the Gaussians ($G$) selected for removal $G_{\text{remove}}$ at a removal iteration are those satisfying:
$G_{\text{remove}}=\{G_{i}\mid d(c_{G_{i}},p^{*})<d_{\text{remove}}\}$ (5)
The removal distance is defined as:
$d_{\text{remove}}=t_{r}\cdot d_{\text{avg}}$ (6)
The removal threshold defaults to $t_{r}=0.5$, a value we determined empirically. This process ensures that the refinement specifically targets meaningful splats and maintains clean, artifact-free refinement.
III-C Gaussian Rasterization
Our rasterization process consists of two key stages, see Fig. 2. First, we build upon semantic GS, in our case GG [8]. This initialization and pre-training phase enables the model to distinguish and categorize scene elements effectively, leveraging semantic information for a strong foundation in subsequent refinements.
In the main phase of our approach, which focuses on POI refinement, we utilize the standard GS rasterizer while employing optimizations to enhance efficiency. Specifically, we accelerate the training process using a fused-SSIM loss alongside sparse Adam optimizer, significantly improving convergence speed [7]. Additionally, we integrate our POI masking mechanism, see Section III-A, allowing for targeted refinement of relevant regions while optimizing a clean scene representation and loss computation. This selective strategy ensures that our approach concentrates on relevant areas.
(a) 1000 iterations
(b) 2000 iterations
(c) 3000 iterations
(d) 4000 iterations
Figure 4: Initialization example on SF_6thAndMission_medium10 from the Neo360 dataset [14]. The rendering iteration is denoted below each column.
TABLE I: Evaluation on NeRDS 360 [14]. We select “car” as POI.
GG [8] (“convex hull”) GG [8] (direct removal) Ours
[…]
6thAndMission_medium
6 12.545 0.547 0.331 2560 20.184 0.724 0.245 2560 27.776 0.959 0.061 418
7 18.858 0.753 0.187 1957 25.732 0.828 0.139 1957 29.790 0.967 0.045 469
10 10.808 0.484 0.366 2061 23.164 0.865 0.162 2061 27.466 0.964 0.058 447
GrantAndCalifornia
1 11.324 0.433 0.298 1848 24.603 0.821 0.164 1848 32.094 0.968 0.052 446
2 10.890 0.433 0.354 2005 22.567 0.758 0.184 2005 31.292 0.960 0.058 446
3 10.341 0.445 0.349 1867 20.030 0.711 0.221 1867 29.822 0.963 0.051 443
VanNessAveAndTurkSt
3 13.459 0.601 0.275 1939 24.469 0.750 0.197 1939 29.010 0.957 0.069 464
5 17.131 0.602 0.271 2160 24.469 0.750 0.197 2160 29.269 0.971 0.047 480
Mean 13.170 0.537 0.304 2050 22.636 0.773 0.191 2050 29.590 0.964 0.055 460
IV EVALUATION
IV-A Metrics
To evaluate the novel view synthesis (NVS) quality of our approach, we report peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and similarity index measure (SSIM). We report the total training time in seconds.
IV-B Implementation Details
All scenes were evaluated on desktop PCs with an RTX3090 with 24 GB VRAM.
For our periodic scene filtering, we use parallelization to calculate $p^{*}$. For this, we use a thread pool with $N$ workers so that several views can be processed simultaneously. To maintain visual fidelity during the refinement process, periodic filtering of the Gaussians based on color similarity is performed every 1000 training iterations.
IV-C Datasets
NeRDS 360 [14]
We use the COLMAP subset from the NeRDS 360 dataset [14]. The scenes are particularly interesting as they show urban landscapes and vehicles, while offering a 360 view of a street scene. The respective scenes with COLMAP ground truth are two from SF_VanNessAveAndTurkSt (3,5), the scenes 1, 2, and 3 from SF_GrantAndCalifornia, and three scenes from 6th and Mission (6, 7, 10). As POI we defined the class “car”.
SCRREAM [13]
For an indoor proof of concept, we evaluated our approach on scenes from the SCRREAM [13]. Each scene contains separate recordings with a variety of objects. For our comparison we use Scene01, Scene02, and Scene03 all in the setting full. We use the class chair as POI.
IV-D Baselines
For the comparison, we selected GG [8], as it is the state-of-the-art in GS separation. Furthermore, as reported by Ye et al.[8], it has the same performance in NVS as vanilla GS. We exclude other approaches like GAGA [12], due to several needed subsequent training steps, and simultaneous localization and mapping (SLAM)-based approaches [33], as they require longer training times.
TABLE II: Comparison of GS after 15k and 30k training iterations. We mask PSNR and SSIM with the GT object mask of the “car”, as GS does not offer per-object extraction.
Iterations PSNR $\uparrow$ SSIM $\uparrow$ TIME $\downarrow$
GS 30000 21.40 0.862 417
Ours 15000 21.73 0.859 285
IV-E Runtime
IV-E1 NeRDS 360
To evaluate the runtime, we measure the overall training time of our approach (CoRe-GS), and the compared baseline (GG with “convex hull” removal, and direct removal), including time for semantic object extraction. As shown in Table I our overall runtime is 460 seconds, while GG, requires 2050 seconds in total. Adding segmentation to the GS rasterization process can enhance quality, but also increases the training time. Our approach only requires a segmentation and a ready-to-explore scene (3k), and we achieve a maximum training time of around 460 seconds in average, where around 120 seconds can be attributed to the initialization training and around 4 seconds to the POI extraction. This shows that our approach reduced the training load enormously.
IV-E2 SCRREAM
Similarly, on the SCRREAM dataset, see Table IV, we require 816 seconds on average while GG takes 2675 seconds. On the SCRREAM dataset, our runtime is more influenced by the number of images used for the POI refinement compared to the NeRDS 360 dataset. The number of images per scene, and per POI is mostly similar on NeRDS 360, while for the non-360 scenes (SCRREAM) the images containing POIs vary more.
(a) No filtering
(b) Filtering
Figure 5: Color-based filtering (cropped to outlier) on SF_GrantAndCalifornia2. Our approach without filtering shows outliers (left), while our filtering leads to a clean background (right).
IV-E3 Comparison to Vanilla Gaussian Splatting
GG and in general semantic GS have prolonged training times due to the extended rasterization process. Since its introduction, the codebase of vanilla GS [7] has seen significant improvements, including a reduction in training time. When comparing this vanilla GS with our approach (CoRe-GS), see Table II, we demonstrate that CoRe-GS, initialized with semantic GS, outperforms vanilla GS in both runtime and PSNR. For this comparison, we used masked PSNR and masked SSIM as traditional GS does not allow the removal of specific scene objects. Our approach achieves a higher PSNR in just half the number of training iterations, while SSIM shows almost the same performance. Furthermore, the overall training time is lower when using our approach.
IV-F Optimizations
CoRe-GS proposes several optimizations for semantic object extraction. While we show the impact on runtime and NVS quality separately on two datasets, we also focused on the joint impact of these. In the following, we evaluate the impact of our chosen number of initializations, as well as the impact of our proposed color-based filtering.
IV-F1 Initialization
For the initialization, we analyzed the trade-off between creating a visually explorable scene and minimizing training time. We evaluated subjective visual quality as shown in Figure 4, and objective metric-based results, see Table III.
The visual quality shows that using a low-quality semantic GS scene for initialization can reduce the final quality within a range of $\pm 0.5$ for PSNR. Although this difference is not high per se, when looking at Figure 4 for a visual scene exploration to select POIs, the benefits of additional iterations become clear. 3000 to 4000 iterations show clearer results. Similarly, in metrics, we can observe that when even using 3000 iterations, combined with our proposed optimizations, a better NVS quality can be achieved. Thus, 3000 iterations provide a positive balance between training time and NVS quality.
IV-F2 Filtering Impact
In Fig. 5, we observe noticeable outliers when color-based filtering is not applied, compared to when it is used. Quantitatively, for the scene SF_GrandAndCalifornia2, the results without filtering are 31.038 PSNR, 0.959 SSIM, and 0.060 LPIPS. With our color-based filtering, PSNR improves by 0.253, SSIM improves by 0.001, and LPIPS improves by 0.002. Although the metric gains are most pronounced for PSNR, the qualitative improvement in Fig. 5 clearly highlights the reduction of visible outliers.
IV-G Novel View Synthesis Quality
IV-G1 NeRDS 360
We compare the overall performance on all datasets with a focus on our speedup compared to GG. As shown in Table I, our approach outperforms GG. When using both settings for GG (“convex hull”, and direct removal). Our approach not only achieves this level of performance but also demonstrates a significantly reduced training time. We outperform GG in all visual metrics. In Figure 3, we provide a direct comparison of filtered outputs between GG and our method. A comparison of CoRe-GS with GG using “convex hull” removal reveals that the “convex hull” method retains numerous outliers (floater) from the original scene. Despite these differences, our approach (CoRe-GS) exhibits no significant drawbacks in terms of NVS quality. The “convex hull” clearly leaves in scene parts, one would like to remove, like the street. While the “direct” removal shows a better performance, it still leaves outliers in the scene, negatively influencing the visual quality.
(a) $G$ convex
(b) $G$ non-convex
(c) Ours
Figure 6: SCRREAM scene01 (images cropped to POI).
TABLE III: Initialization iterations (1k-4k). We evaluate after 7k/30k final iterations on 6thAndMission_medium10 [14].
initial 7k final iterations 30k final iterations
[…]
1000 26.628 0.958 114 27.107 0.962 380
2000 27.054 0.961 140 27.542 0.965 391
3000 27.105 0.962 180 27.668 0.965 447
4000 27.117 0.962 224 27.553 0.964 480
IV-G2 SCRREAM
To evaluate our approach in multiple settings, we also compared it on an indoor dataset. As shown in Figure 6, the same issues for GG as for the NeRDS dataset can usually be observed. The SCRREAM dataset [13] (real-world, indoor) consists of diverse scenes with highly accurate ground truth for semantic segmentation. As shown in Table IV, our approach outperforms GG in two scenes. In the second scene, a lower visual quality can be observed for both approaches, with GG showing a better result, in this case. Here, CoRe-GS is challenged and leaves in floaters as well, resulting in the low visual quality. Although this problem can be mitigated using a different $t_{r}$ value, we choose to keep the same hyperparameters for all scenes, to retain a fair comparison. However, a potential optimization can be seen in Fig. 7. Despite scene02, overall we outperform GG with an improvement in PSNR of 0.318, in SSIM of 0.046, and in LPIPS of 0.064.
Figure 7: Limitation. Exemplary image from scene02 of SCRREAM illustrating a failure case of CoRe-GS. The masked GT (left) is a rendering result from Table IV, and CoRe-GS using $t_{r}=1.0$ (right), instead of $t_{r}=0.5$ (center).
TABLE IV: Evaluation on SCRREAM [13]. We use the scenes 01/02/02 (_full_00) and select “chair” as POI.
$G$ (direct) [8] Ours
[…]
01 45.521 0.973 0.018 2614 45.799 0.994 0.014 709
02 21.663 0.829 0.056 2766 15.207 0.920 0.262 1083
03 45.521 0.973 0.018 2645 52.653 0.997 0.009 656
Mean 37.568 0.925 0.031 2675 37.886 0.970 0.095 816
V DISCUSSION
We propose CoRe-GS, a novel method to refine POI for GS. Complete reconstruction of a scene is computationally intensive and often unnecessary when only certain POIs are relevant for timely decision-making. Our approach solves this problem with a hybrid solution: while a segmentation-ready scene can be examined early on, high-quality refinements are tailored to POIs, reducing the total training time to 25% compared to full semantic GS training. Our approach leverages semantic GS [8] in the first stage to initialize the POI refinement step. The semantic GS module is interchangeable and independent from the POI refinement, fostering modularity and scalability. Our color-based filtering and POI selection enable an efficient training cycle, ensuring that only data from the target object contributes to the scene optimization. This results in a more focused and faster training process by ignoring irrelevant background information.
As shown in Table I, the runtime of GG is notably higher compared to our approach. While semantic information is necessary for selecting POIs after obtaining the coarse, segmentation-ready scene representation, our method provides a substantial speedup and extracts more visually appealing POIs (Figure 3).
Our main motivation lies in situations where operators can select POIs from drone footage, which is why we focus mainly on outdoor data within the NeRDS 360 dataset. Experiments on the SCRREAM dataset validate the robustness of our method across diverse environments and highlight that further improvements may increase this robustness in indoor settings. Although the relative advantage over the baselines is smaller than on the outdoor NeRDS 360 scenes, our method still produces strong results. This translation to datasets outside of our intended application use case shows the scalability of our approach.
Additionally, the POI refinement gives CoRe-GS an advantage in occluded areas, see Fig. 3. In the complete scene, a pole or street light can occlude the POI from different NVS perspectives. Selecting a POI and extracting offers a more complete understanding of the object’s geometry and structure from the same perspective. This is not achievable with GG and “convex hull” removal. While direct removal allows it, outliers still negatively impact the result compared to our outlier-free visualization.
VI LIMITATIONS
CoRe-GS benefits from high-quality segmentation. Its performance is therefore constrained by the segmentation quality. This limitation is not unique to our approach [4]. Previous work [4] has shown that the accuracy of segmentation quality has a critical influence onNVS quality and downstream tasks.
VII CONCLUSION
We introduced CoRe-GS, a POI refinement for GS that selectively enhances segmentation-based POIs critical for scene exploration and operator decision making. By focusing computation on targeted scene areas, our approach produces higher-fidelity POIs without the overhead of refining the entire scene representation. Our evaluation on two benchmarks demonstrates clear advantages over both standard GS and GG in two object extraction settings. Unlike standard GS, which lacks POI support, our method achieves targeted visual refinement and, compared to GG, it reduces training time while preserving segmentation accuracy and rendering quality. These results highlight our approach’s efficiency and visual quality, while its modular design offers potential for future applications in more complex and specialized environments and mobile reconstruction scenarios.
ACKNOWLEDGMENTS
This work has been partially supported by the Technical University of Munich, Munich, Germany, within its MIRMI seed fund scheme.